OpenStudy (anonymous):

Write a system of equations to solve the following problem. Let c be the number of child tickets and a be the number of adult tickets. Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket were sold?

OpenStudy (anonymous):

Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket were sold? Answer choice: 25 child and 8 adult

hero (hero):

@ILuvHomework, the directions say to create "a system of equations" before solving the problem. Please do not give answers. Explain, Guide and tutor. Thank you.

OpenStudy (anonymous):

@Hero ? Do you think you could help me to understand? :)

hero (hero):

@Amber1997, how many tickets were sold in total?

OpenStudy (anonymous):

33

hero (hero):

Very good. And the kinds of tickets sold are either childrens tickets (c) or adult tickets (a) Right?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Okay I think I got it.

hero (hero):

So basically the sum of children's tickets and adult tickets sold will be equal to the 33 total tickets sold. Do you have an idea of how we represent that as an equation?

OpenStudy (anonymous):

3c + 5a = 115 c + a = 33 Is this equation right?

hero (hero):

Okay. Just to be sure you understand, can you describe what the equation 3c + 5a = 115 means in words?

OpenStudy (anonymous):

three dollars multiplied by the number of children attending, plus five dollars times the number of adults attending, equals one-hundred-fifteen dollars

hero (hero):

One equation represents the total amount of tickets sold. The other equation represents the total amount of money collected based on the cost of each ticket and the amount of tickets sold.

hero (hero):

Okay, very well. My next question is, do you know how to go about solving the system of equations?

OpenStudy (anonymous):

Yes, that I understand! :) Thank you for your help @Hero , I really appreciate it!

hero (hero):

@Amber1997, have you already solved the system? If so, would you mind showing the work you did to solve it?

OpenStudy (anonymous):

@Hero I have a question. Is the answer 5c + 3a = 115 and c + a = 33? Sorry if that is direct. I was doing some homework and that is what I got as the answer. But when reading what you have said before, it makes me a little confused. I do not fully understand what you are saying. Do you think you could help me?